# The Open Journal of Astrophysics

### v1

@@ 3.3.2 Halo cross correlations
Figure 9: The halo cross correlation error for the SR (dotted lines) and Emulator (solid lines) models for different halo mass regimes (panels from left to right). The color of each curve corresponds to its redshift according to the color bar. The shaded area shows the $1\sigma$ standard deviation from all test sets. The vertical dashed gray lines indicate the Nyquist frequency.
Figure 9 shows the halo stochasticity error. We find that the emulator achieves smaller errors than the SR model across all mass ranges. In the mass range greater than $10^{13}\,M_{\odot}$, the improvements made by Emulator are less significant compared to the enhancements within the mass range of $[10^{11},10^{13}]\,M_{\odot}$. This is because our SR model, being conditioned on LR simulations, already constrains the positions and masses of massive halos. The emulator fine-tunes these positions and masses to more closely match those in the HR field test set. Since the SR model conditions its predictions on LR simulations, it is expected that the correlation between the SR field and the HR field diminishes from $10^{13}\,M_{\odot}$ down to $10^{11}\,M_{\odot}$. However, by adjusting halos based on the HRICs, our emulator achieves much higher accuracy within this mass range. In the mass regime below $10^{11}\,M_{\odot}$, the HRICs lack sufficient information on less massive halos, resulting in both the SR field and emulator output failing to predict halos that match the positions and masses of those in the HR field.
@@ 3.3.3 Merger histories
Given the emulator’s success in reproducing the positions and masses of individual halos, it is interesting to examine whether the merger histories of these halos are also consistent the HR run. To do this, we use the merger tree algorithm built into the cosmological $N$-body code Gadget-4 (Springel et al., 2021), called Hierarchical Bound Tracing (HBT). It uses information from earlier snapshots to help derive the latest halo catalogue. Starting from high redshift, the algorithm identifies main halos at their formation and follows the particles within these halos through subsequent snapshots. This process generates a merger tree for the main halos at the first stage, which is then extended to include subhalos.
Figure 10 shows a direct visual comparison of merger trees for the same halo selected from the SR field, Emulator output and the HR field (from bottom to top, redshift changes from $z=1$ to $z=0$). This halo, with mass $\sim 10^{14}\,M_{\odot}$ was selected based on and mass relative to the HR run. The size of each tree node is scaled according to halo mass, with red colors indicating the main progenitors. Although the SR merger tree looks authentic, with mass fluctuations and mergers, tthe redshifts when the mergers happen and their numbers are very different in detail from the counterpart HR halo. In contrast, the Emulator’s merger tree more accurately predicts the redshift and masses involved in the main merger event. The emulator output also features more small-mass halos, i.e., more substructures than the SR field, and places them in relatively correct locations as a function of time.
@@
To further validate the actual merger history, such as mass accretion history measured from the SR, Emulator and HR fields, we select halos within different mass ranges from the SR and emulator fields and calculate the mass error relative to the corresponding HR halos from $z=1$ to $z=0$. First, we split all the halos in the HR run at $z=0$ into four different mass regimes, then track the mass history of each halo along its main progenitor branch from $z=1$ to $z=0$. To demonstrate the accuracy of our emulation and SR model, we calculate the mass error using the following equation:
@@ 4 Discussion
However, there remain some discrepancies in agreement with the HR runs. Both the SR model and the Emulator currently exhibit a lack of substructures within very massive halos, particularly at the higher mass end, as seen in Figure 7 and the visual comparisons in Section 3.1. Most of the missing halos are of low mass, suggesting that increasing the resolution of the Eulerian fields, used to provide direct information about substructures to the discriminator, could improve this. At tte current spatial resolution ($\approx 0.2h^{-1}{\rm Mpc}$), the discriminator struggles to identify substructures within massive halos that require higher contrast to resolve.
@@ 5 Summary and Conclusions
We have developed and evaluated an emulator based on neural networks that produces super-resolved cosmological $N$-body simulation outputs. The inputs are a low-resolution $N$-body simulation and the initial conditions from a high-resolution simulation. The Emulator processes the outputs from the SR model of Zhang et al. 2023, which is based on an LR simulation. By integrating the initial conditions of particles (including initial displacement and velocity), the emulator predicts the adjustments to the SR model’s 6D phase space output. These adjustments ensure that emulator’s output closely aligns with the HR $N$-body simulation using the same initial conditions.
We conducted detailed validation of our Emulator by comparing various statistics calculated from test sets. In Section 3.1, we presented a visual comparison between the SR model, the Emulator and the corresponding HR simulation. These comparisons demonstrate that on large scales, the SR, emulator and HR fields appear identical. While SR field presents visually convincing small-scale structures, it differs in detail from the HR simulations. However, the Emulator output successfully aligns with the HR simulation even on scales well below the resolution of the LR simulation.
In Section 3.2, we examined the outputs in Fourier space and found that the emulator output correlates much more closely with the HR field than SR output, from $k=1h{\rm Mpc}^{-1}$ down to the Nyquist frequency. The Emulator’s accuracy represents a significant improvement over the SR field. To evaluate the halo population and substructures, we compute the halo and subhalo mass functions, as well as mean occupation number, in Section 3.3. Both the SR and Emulator host halo mass abundance have good statistical agreement with HR, showing our emulator does not diminish the previous results from SR field (Zhang et al. 2023). At the same time, the subhalo mass function shows that the Emulator output improves significantly over the SR field, adjusting the SR field to the HR field and improving the prediction of small-scale structures. However, a deficit in the subhalo abundance for halos of mass $[10^{13},10^{14}]\,M_{\odot}$ indicates that current model still struggles to predict a sufficient number of substructures within large mass halos, likely due to the limitations of the Eulerian field resolution.
@@
Our findings demonstrate the emulator’s capability to adjust the SR field output to match the corresponding HR field by utilizing the initial conditions. Our Emulator serves as a promising foundation for emulating $N$-body simulations under different constraints, suggesting the feasibility of generally emulating outputs for simulations with varying cosmological parameters across a wide range of redshifts. The integration of the SR model with the emulator framework enables the rapid and accurate generation of results across various redshifts, providing an effective tool which can compete with other approximate methods such as those of Feng et al. (2016) and Scoccimarro & Sheth (2002), but promising more accuracy and the potential inclusion of more physics. The emulator approach allows for the generation of mock halo catalogs, as well as other non-linear statistics (such as computing power spectrum covariance matrices), and should become a useful tool for facilitating the constraint of cosmological parameters from cosmological surveys.
@@ Appendix A Results from Training the Emulator Without Discriminator Guidance
We now present the results of training the same Emulator without discriminator guidance. In this context, "without discriminator guidance" means training the neural network using the loss function in Equation 6, but without the third adversarial term.
Figure 13 shows the result from the Emulator without discriminator guidance (hereafter referred to as Emulator w/o DG). The first row shows the whole $(100h^{-1}{\rm Mpc})^{3}$ volume, centered on the same halo as in Figure 2 and Figure 3 , to show the large-scale structures, the second row and the third row shows a sub-volume of $(20h^{-1}{\rm Mpc})^{3}$ and $(11h^{-1}{\rm Mpc})^{3}$, to show the different detailed structures and halo distributions. Comparing with Figure 2 and Figure 3, we can see that the Emulator w/o DG still appears authentic on large scales, similar to our Emulator+DG model. However, as shown in the third column, on smaller scales, the Emulator w/o DG tends to create a much smoother density field, resulting in a lack of substructures.
Figure 14 shows the results of stochasticity, calculated using the same equations as in Section 3.2. Comparing these results with those in Figure 4, we can see that the Emulator w/o DG performs slightly better in the density field and produces comparable results in displacement and momentum.
The host-halo mass function, sub-halo mass function, and mean occupation number for the Emulator w/o DG’s output are shown in Figure 15. Both the host-halo and sub-halo mass functions show a deficit in the lower mass region, indicating that training without discriminator guidance (DG) lacks the detailed structures observed on smaller scales in the Emulator+DG model. Additionally, the mean occupation number reveals that training without DG results in fewer low-mass subhalos within the range of $[10^{11},10^{13}]\,M_{\odot}$.
@@
Figure 13: 3D visualization of the Emulator w/o DG density fields with boxsize $100h^{-1}{\rm Mpc}$ (first column), $20h^{-1}{\rm Mpc}$ (second column), $11h^{-1}{\rm Mpc}$ (third column) at $z=1$ and $z=0$.
@@
Figure 15: The host halo, subhalo mass function and mean occupation number at redshifts $z=1,0.8,0.5,0.2,0$ measured from the Emulator w/o DG output. The color of each curve denotes the redshift of simulation in test set. The shaded areas indicate $1\sigma$ standard deviation from all test sets. The upper panel shows halo mass functions and lower panel shows fractional errors

### v2

@@ 3.3.2 Halo cross correlations
Figure 10: The halo cross correlation error for the SR (dotted lines) and Emulator (solid lines) models for different halo mass regimes (panels from left to right). The color of each curve corresponds to its redshift according to the color bar. The shaded area shows the $1\sigma$ standard deviation from all test sets. The vertical dashed gray lines indicate the HR Nyquist frequency and the dotted gray lines indicate the LR Nyquist frequency.
Figure 10 shows the halo stochasticity error. We find that the emulator achieves smaller errors than the SR model across all mass ranges. In the mass range greater than $10^{13}\,M_{\odot}$, the improvements made by Emulator are less significant compared to the enhancements within the mass range of $[10^{11},10^{13}]\,M_{\odot}$. This is because our SR model, being conditioned on LR simulations, already constrains the positions and masses of massive halos. The emulator fine-tunes these positions and masses to better match those in the HR field test set. Since the SR model conditions its predictions on LR simulations, it is expected that the correlation between the SR field and the HR field diminishes from $10^{13}\,M_{\odot}$ down to $10^{11}\,M_{\odot}$. However, by adjusting halos based on the HRICs, our emulator achieves significantly higher accuracy within this mass range. In the mass regime below $10^{11}\,M_{\odot}$, the HRICs lack sufficient information on less massive halos, resulting in both the SR field and emulator output failing to predict halos that match the positions and masses of those in the HR field.
@@ 3.3.3 Merger histories
Given that the Emualtor is capable of reproducing the positions and masses of individual halos, it is interesting to examine whether the merger histories of these halos are also consistent with the HR run. To do this, we use the merger tree algorithm built into the cosmological $N$-body code Gadget-4 (Springel et al., 2021), called Hierarchical Bound Tracing (HBT). It uses information from earlier snapshots to help derive the latest halo catalog. Starting from high redshift, the algorithm identifies main halos at their formation and follows the particles within these halos through subsequent snapshots. This process generates a merger tree for the main halos at the first stage, which is then extended to include subhalos.
Figure 11 shows a direct visual comparison of merger trees for the same halo selected from the SR field, Emulator output and the HR field (from bottom to top, redshift changes from $z=1$ to $z=0$). This halo, with mass $\sim 10^{14}\,M_{\odot}$ was selected based on the mass relative to the HR run. The size of each tree node is scaled according to halo mass, with red colors indicating the main progenitors. Although the SR merger tree looks authentic, with mass fluctuations and mergers, the redshifts when the mergers happen and their numbers are very different in detail from the counterpart HR halo. In contrast, the Emulator’s merger tree provides more accurate predictions of the redshift and masses involved in the main merger event. Additionally, the Emulator output includes a greater number of small-mass halos, i.e., more substructures compared to the SR field, and positions them in relatively accurate locations as a function of time..
@@
To further validate the actual merger history, such as mass accretion history measured from the SR, Emulator and HR fields, we select halos within different mass ranges from the SR and emulator fields and calculate the mass error relative to the corresponding HR halos from $z=1$ to $z=0$. First, we split all the halos in the HR run at $z=0$ into three different mass regimes, then track the mass history of each halo along its main progenitor branch from $z=1$ to $z=0$. To demonstrate the accuracy of our emulation and SR model, we calculate the mass error using the following equation:
@@ 4 Discussion
However, there are some discrepancies in agreement with the HR runs. Both the SR model and the Emulator currently exhibit a lack of substructures within very massive halos, particularly at the higher mass end, as seen in Figure 8 and the visual comparisons in Section 3.1. Most of the missing halos are of low mass, suggesting that increasing the resolution of the Eulerian fields, used to provide direct information about substructures to the discriminator, could improve this. At current spatial resolution, the discriminator struggles to identify substructures within massive halos that require higher contrast to resolve.
@@ 5 Summary and Conclusions
We have developed and evaluated an emulator based on neural networks that produces super-resolved cosmological $N$-body simulation outputs. The inputs to the Emulator include a low-resolution $N$-body simulation and the initial conditions from a high-resolution simulation. The Emulator refines the outputs of the SR model presented in Zhang et al. 2023, which is based on the LR simulation. By integrating the initial conditions of particles (including initial displacement and velocity), the Emulator predicts the adjustments to the SR model’s 6D phase space output. These adjustments ensure that the emulator’s output closely aligns with the HR $N$-body simulation using the same initial conditions.
We conduct detailed validation of our Emulator by comparing various statistics calculated from test sets. In Section 3.1, we present a visual comparison between the SR model, the Emulator and the corresponding HR simulation. These comparisons demonstrate that on large scales, the SR, Emulator and HR fields appear identical. While SR field presents visually convincing small-scale structures, it differs in detail from the HR simulations. However, the Emulator output successfully aligns with the HR simulation even on scales well below the resolution of the LR simulation. We also compared the error distributions between the SR model and the Emulator. The Emulator significantly reduces particle displacement and velocity errors compared to the SR model.
In Section 3.2, we analyze the outputs in Fourier space and found that the emulator output correlates much more closely with the HR field than SR output, from $k=1h{\rm Mpc}^{-1}$ down to the Nyquist frequency. The Emulator’s accuracy represents a significant improvement over the SR field. To evaluate the halo population and substructures, we compute the halo and subhalo mass functions, as well as mean occupation number, in Section 3.3. Both the SR and Emulator host halo mass abundance have good statistical agreement with HR, showing our Emulator does not diminish the previous results from SR field (Zhang et al. 2023). At the same time, the subhalo mass function shows that the Emulator output improves significantly over the SR field, adjusting the SR field to the HR field and improving the prediction of small-scale structures. However, a deficit in the subhalo abundance for halos of mass $[10^{13},10^{14}]\,M_{\odot}$ indicates that current model still struggles to predict a sufficient number of substructures within large mass halos, likely due to the limitations of the Eulerian field resolution.
@@
Our results demonstrate the Emulator’s ability to adjust the SR field output to closely match the corresponding HR field by leveraging the initial conditions. The Emulator represents a promising foundation for emulating $N$-body simulations under different constraints, suggesting the feasibility of generally emulating outputs for simulations with varying cosmological parameters across a wide range of redshifts. The integration of the SR model with the Emulator framework enables the rapid and accurate generation of results across various redshifts, providing an effective tool which can compete with other approximate methods such as those of Feng et al. (2016) and Scoccimarro & Sheth (2002), while offering greater accuracy and the potential inclusion of more physics. The Emulator approach facilitates the generation of mock halo catalogs and the computation of non-linear statistics, such as power spectrum covariance matrices, positioning it as a valuable tool for constraining cosmological parameters from upcoming cosmological surveys.
@@ Appendix A Results from Training the Emulator Without Discriminator Guidance
We present the results of training the same Emulator without discriminator guidance. In this context, "without discriminator guidance" means training the neural network using the loss function in Equation 6, but without the third adversarial term.
Figure 14 shows the result from Emulator w/o DG model. The first column shows the whole $(100h^{-1}{\rm Mpc})^{3}$ volume, centered on the same halo as in Figure 2 and Figure 3, to illustrate the large-scale structures, the second and the third columns show a sub-volume of $(20h^{-1}{\rm Mpc})^{3}$ and $(11h^{-1}{\rm Mpc})^{3}$, to show the different detailed structures and halo distributions. Comparing with Figure 2 and Figure 3, we can see that the Emulator w/o DG still appears authentic on large scales, similar to our Emulator+DG model. However, as shown in the third column, on smaller scales, the Emulator w/o DG model produces a significantly smoother density field, leading to a noticeable lack of substructures.
Figure 15 shows the results of stochasticity, calculated using the same equations as in Section 3.2. We can see that the Emulator w/o DG performs slightly better in the density field and produces comparable results in displacement and momentum.
The halo mass function, subhalo mass function, and mean occupation number for the Emulator w/o DG’s output are shown in Figure 16. Both the halo and sub-halo mass functions show a deficit in the lower mass region, indicating that training without discriminator guidance (DG) lacks the detailed structures observed on smaller scales in the Emulator+DG model. Additionally, the mean occupation number reveals that training without DG results in fewer low-mass subhalos within the range of $[10^{11},10^{13}]\,M_{\odot}$.
@@
Figure 14: 3D visualization of the Emulator w/o DG density fields with boxsize $100h^{-1}{\rm Mpc}$ (first column), $20h^{-1}{\rm Mpc}$ (second column), $11h^{-1}{\rm Mpc}$ (third column) at $z=1$ and $z=0$. The first and third rows correspond to the Emulator w/o DG (without discriminator guidance), while the second and fourth rows represent the standard Emulator.
@@
Figure 16: The halo, subhalo mass function and mean occupation number at redshifts $z=1,0.8,0.5,0.2,0$ measured from the Emulator w/o DG output. The color of each curve denotes the redshift of simulation in test set. The shaded areas indicate $1\sigma$ standard deviation from all test sets. The upper panel shows halo mass functions and lower panel shows fractional errors